# DarkGS: Learning Neural Illumination and 3D Gaussians Relighting for Robotic Exploration in the Dark

### v1

@@ Abstract
Humans have the remarkable ability to construct consistent mental models of an environment, even under limited or varying levels of illumination. We wish to endow robots with this same capability. In this paper, we tackle the challenge of constructing a photorealistic scene representation under poorly illuminated conditions and with a moving light source. We approach the task of modeling illumination as a learning problem, and utilize the developed illumination model to aid in scene reconstruction. We introduce an innovative framework that uses a data-driven approach, Neural Light Simulators (NeLiS), to model and calibrate the camera-light system. Furthermore, we present DarkGS, a method that applies NeLiS to create a relightable 3D Gaussian scene model capable of real-time, photorealistic rendering from novel viewpoints. We show the applicability and robustness of our proposed simulator and system in a variety of real-world environments. Code released at https://github.com/tyz1030/neuralight.git
@@ I INTRODUCTION
Figure 1: Robotic imaging systems working in the dark consist of cameras and light sources. Examples as shown in (a): Carla Simulator [1], Team CoStar in SubT Challenge [2] and HoloOcean underwater robot simulator [3]. In this work, we propose a pipeline that calibrates the camera-light system which helps photorealistic scene reconstruction and relighting from images collected in the dark.
Scene reconstruction, or the capability to create accurate internal representations of the environment, is vital for robots operating in unknown environments. Previous vision-only approaches largely rely on identifying common feature points over a set of multiple-view images, and then minimizing a reprojection error [4]. Such procedures like SfM (SfM) or SLAM (SLAM) also estimate the camera poses of the images. Based on these camera poses, NeRF (NeRF[5] is capable of achieving photorealistic scene reconstruction by optimizing a photometric loss between the representation and the images. However, while achieving huge success in the graphics community, the transition of objective function from reprojection error to photometric loss has raised a new challenge to the robotics community: Can we still build consistent scene representations with a moving light source on the robot platform?
In this work, we identify illumination-inconsistency as the main challenge in building a photorealistic scene representation from images taken with a moving light source. That is, as the robot operates in the environment, the effect of moving light source results in images captured of the same region appearing visually different. Handcrafted feature descriptors that serve as backbone in SfM and SLAM [6, 7, 8] are designed to be invariant to changes in pixel intensities. Such changes also have limited effects on the optimization pipeline, which uses reprojection errors as objective functions. Hence, the classic SfM methods are generally robust against illumination-inconsistency. However, on the way towards photorealistic scene reconstruction and switching to optimization of the photometric loss, we find that state-of-the-art NeRF methods [5, 9] fail on the images taken in the dark with a moving light source. Although variants such as RawNeRF [10] and Relightable 3DG [11] claim to be able to handle dark images in RAW format and model the reflection properties of the scene, they still fail to handle the varying illumination.
@@
    DarkGS (DarkGS), a variant of the 3DGS (3DGS) model that builds photorealistic scene representations under poorly illuminated conditions and relights the scene with global illumination, based on COLMAP [12] and NeLiS results.
@@ II-A Extroceptive Sensor Calibration on Robots
Autonomous robots are usually equipped with perceptual sensors such as cameras and LiDARs. Taking camera calibration as an example, a calibration target with AprilTag [13] or checkerboard pattern is often used. With observations of the target from different perspectives, the focal length, center of projection, distortion coefficient, and pose of the camera can be estimated [14]. Similar target-based approaches have been used to calibrate LiDARs [15], radars [16] and acoustic sensors [17] for downstream sensor fusion tasks [18]. Analogous to sensor calibrations mentioned above, in this paper, we propose to use a target consisting of AprilTags [13] and blank space for calibrating the light in a camera-light system, including estimating the transformation between camera’s and light’s coordinate system, RID (RID) and light fall-off curve of the light source.
@@ II-B Light Calibration
Existing methods use various kinds of customized calibration targets: [19][20] propose to use a target of AprilTags and pins to infer the position of point light source from the shadows of the pins. Alternatively,  [21] [22] propose to use a Lambertian sphere and estimate the light source parameters by learning to reconstruct the sphere. Furthermore, [23] shows that the location of the light source can also be recovered from the shadow of a sphere. With a Lambertian plane and customized markers,  [24] shows that it is possible to calibrate a point light source with a single image. [25] shows the most relevance to our work, which calibrates the pose of a light source and the metric scale given the RID curve. However, the above work either assumes an oversimplified model, e.g. a point light source with inverse square light falloff, in a pure dark environment, or that the RID is given, which makes the method less applicable. In this work, we use a minimum calibration target design that consists of only AprilTags [13] and a Lambertian plane similar to [25]. With NeLiS we show that modeling and learning RID, light falloff and ambient light together with camera-light pose can improve the quality of image reconstruction and make our calibration method more feasible for real-world robot deployments.
@@ II-C NeRF and 3D Gaussian Splatting
Based on the success of SfM [4], NeRF learns to represent scenes using a continuous function, e.g. a MLP (MLP), and can achieve photorealistic novel-view image synthesis. Popular variations of NeRF such as Mip-NeRF[26] and Instant-NGP [27] have improved NeRF by rendering accuracy and speed. Recent innovation introduced as 3DGS [9] has revolutionized this field by replacing the neural network backbone with spatial Gaussians, allowing rendering speed of 100+ fps while maintaining the accuracy and differentiability of the system. For our robotic setup, the problem of varying illumination is attempted by [28, 29] which model the physical property of the scene, e.g. roughness, albedo, and normal, instead of the radiance as a constant. However, these works oversimplify the light model as a Lambertian point light source, co-centered with the camera, making it not applicable on many real robotic platforms. Gaussianshader [30] and Relightable 3DG [11] introduce physical properties into the 3DGS framework but model illumination as a constant that does not change frame-by-frame. According to our experiments, none of the above mentioned methods handles the illumination-inconsistancy issue on a real-robot imaging system. In our proposed DarkGS, by modeling the physical property of the scene and taking advantage of NeLiS, we can not only build a consistent 3DGS from poorly illuminated images, but also relight the scene with global illumination.
@@ II-D Photometric Stereo
Photometric stereo refers to the estimation of the shape of an object with multiple images taken from a single viewpoint with different lighting orientations [31]. Early approaches confine this problem to known light source and Lambertian surfaces with uniform albedo. Further work has allowed spatially varying albedo [32] and unknown light [33]. However, our robotic setup differs from this problem by the fact that our camera and light source always move together as a rigid body. And we choose to tackle this problem within the framework of 3DGS [9] which allows us to synthesize novel views with photorealistic quality.
@@ III METHODOLOGY
Our proposed workflow for building NeLiS and DarkGS is shown in Fig. 2. We build NeLiS by calibrating the camera-light system at a planar calibration target. By accounting for the calibration image sequence and hand measurement of the light pose, an in-the-loop human operator can initialize the light parameters using our intuitive GUI. Since the calibration target can be localized by AprilTags [13], a shading model can render the image with the given parameters to approximate the calibration image. Then, we run the optimization to minimizes the photometric error of the calibration image sequence and rendered image. After obtaining a calibrated NeLiS, the robot can be deployed in a dark environment, using the collected images to build a DarkGS, finetune the model, and relight the scene.
@@ III-A Shading Model for NeLiS
Given the true size of the calibration target, we can apply PnP (PnP[34] to extract $R_{c}^{w}\in SO(3)$ and $t_{c}^{w}\in\mathbb{R}^{3}$ which transform points from the camera coordinate to the world coordinate. The position of the light source is then given by $\mathbf{o}_{l}=R_{c}^{w}t_{l}^{c}+t_{c}^{w}\in\mathbb{R}^{3}$. Since we align the $z$ axis of the light coordinate with the centerline of the light, the direction of the centerline can be denoted by $\boldsymbol{\omega}_{l}=R_{c}^{w}R_{l}^{c}[0,0,1]^{\top}\in\mathbb{R}^{3}$. Both $\mathbf{o}_{l}$ and $\boldsymbol{\omega}_{l}$ are in the world coordinate frame.
@@ III-A1 RID
RID is commonly modeled as a function of the angle between the centerline $\boldsymbol{\omega}_{l}$ of the light and light ray $\boldsymbol{\omega}_{\mathbf{x}}=\mathbf{x}-\mathbf{o}_{l}$. Previous work assume that RID is given [25, 24, 22]. However, this assumption may not hold for in-the-wild robot deployment. Instead, we remove this dependency by learning a neural RID from calibration data:
@@ III-A2 Light Falloff Curve
Inverse square law is widely used to model light falloff, based on the assumption of a point light source. However, when objects are closer to the light source, the inverse square law starts to fail. We choose to model the light falloff in the form of a Lorentzian function of the distance $\lVert\boldsymbol{\omega}_{\mathbf{x}}\rVert_{2}$, as suggested by [35]:
@@
Instead of estimating $\tau$ from hand measurement of the light surface [35], we designate it to be a learnable parameter.
@@ III-A4 BRDF
Although there are BRDF (BRDF) models capable of handling complex reflection effects, they exceed the scope of this paper. We opt for the Lambertian reflection model, eliminating the need to optimize any parameters in our BRDF. We use $f_{r}(\boldsymbol{\omega}_{\mathbf{x}},\mathbf{n})=\text{max}(\boldsymbol{\omega}_{\mathbf{x}}\cdot\mathbf{n},0)$ as our BRDF, giving the rendering equation:
$\hat{L}_{\mathbf{x}}=I_{\mathbf{x}}f_{r}(\boldsymbol{\omega}_{\mathbf{x}},\mathbf{n})\mathbf{c}$ (4)
@@ III-B Building DarkGS
$\hat{L}_{u,v}=\sum_{i\in\mathcal{N}}I_{i}f_{r}(\boldsymbol{\omega}_{i},\mathbf{n}_{i})\mathbf{c}_{i}\alpha_{i}\prod_{j}^{i-1}(1-\alpha_{j})$ (6)
@@ III-B1 Scale Recovery
The framework of 3DGS and its variants are often based on monocular SfM solutions such as COLMAP [4]. While calibration using AprilTags recovers poses in true metric scale with NeLiS, monocular SfM only gives up-to-scale poses for building 3D Gaussians. Here, we introduce a scaling factor $s>0$ as a learnable parameter so that we can obtain the positions with scale $\mathbf{p}_{i}^{\prime}=s\mathbf{p}_{i}$. With captured pixel intensity $L_{u,v}$, the 3D Gaussian can be built by solving the follow optimization problem:
@@ IV-A Experiments Setup
Our imaging system consists of a FLIR machine vision camera (Model Firefly S) and a light source, together mounted on a rigid structure on top of our legged robot, as shown in Fig. 4. The baseline between the camera and the light source is approximately 32 cm, measured by hand. Although the ray-tracing problem can be simplified by placing the light source and camera close enough and modeling them as co-centered [28], in many applications a long baseline between them is a preferred design worth studying. For example, deep-sea robots usually have a long baseline between light and camera to avoid massive light backscattering from the water.
We experimented with three different light sources: a flashlight, a diving light, and a flood light. To build NeLiS, we take $\sim 40$ calibration images from different ranges and perspectives for each light configuration and use $25\%$ of the images as a testing set to evaluate the NeLiS performance. For building DarkGS, we take $50\sim 150$ images for each scene. All images are streamed in RAW format without any tone mapping or white balancing. Parameters optimized in both NeLiS and DarkGS use Adam optimizer [36] with LieTorch [37] integration.
@@ IV-B Can existing 3DGS methods do the job?
We first investigated whether existing 3DGS methods can reconstruct the scene using the images we collected with our robotic setup in the dark. We experimented with: 1. vanilla 3DGS [9] which models the scene with constant radiance 2. RawNeRF [10] which is developed to reconstruct scene from RAW HDR images 3. Relightable 3DGS [11] which models the shading and environmental light map. To make a fair comparison, we apply gamma tone mapping to images for vanilla 3DGS [9] and Relightable 3DGS [11] which brightens the image and smooth the discrepancy between illuminated and under-illuminated areas. We also replace RawNeRF’s backbone MLP-based renderer [10] with Gaussian render [9].
As shown in Fig. 5, we found that none of the existing methods is able to build valid scene reconstructions. We see an excessive amount of artifacts in the center area of the synthesized image. The key reason is that the capability to model varying illumination is missing from existing methods.
Figure 5: None of the existing methods can solve the problem: Results of Vanilla 3DGS [9], RawNeRF [10] Relightable 3DGS [11] show heavy artifacts and fail to converge. The key reason for the failures is that the existing method does not model the illumination change.
@@
Figure 6: (a) Light sources on real robots have various RID patterns; (b) Real world measurements of light falloff show that the inverse square law is insufficient to model any of our light sources, but Lorentzian functions [35] with learnable parameter $\tau$ fit them well.

### v2

@@ Abstract
Humans have the remarkable ability to construct consistent mental models of an environment, even under limited or varying levels of illumination. We wish to endow robots with this same capability. In this paper, we tackle the challenge of constructing a photorealistic scene representation under poorly illuminated conditions and with a moving light source. We approach the task of modeling illumination as a learning problem, and utilize the developed illumination model to aid in scene reconstruction. We introduce an innovative framework that uses a data-driven approach, Neural Light Simulators (NeLiS), to model and calibrate the camera-light system. Furthermore, we present DarkGS, a method that applies NeLiS to create a relightable 3D Gaussian scene model capable of real-time, photorealistic rendering from novel viewpoints. We show the applicability and robustness of our proposed simulator and system in a variety of real-world environments. Code released at https://tyz1030.github.io/proj/darkgs.html
@@ I INTRODUCTION
Figure 1: Robotic imaging systems working in the dark consist of cameras and light sources. Examples as shown in (a): Carla Simulator [carla], Team CoStar in SubT Challenge [costarnebula] and HoloOcean underwater robot simulator [holoocean]. In this work, we propose a pipeline that calibrates the camera-light system which helps photorealistic scene reconstruction and relighting from images collected in the dark.
Scene reconstruction, or the capability to create accurate internal representations of the environment, is vital for robots operating in unknown environments. Previous vision-only approaches largely rely on identifying common feature points over a set of multiple-view images, and then minimizing a reprojection error [buildrome]. Such procedures like SfM (SfM) or SLAM (SLAM) also estimate the camera poses of the images. Based on these camera poses, NeRF (NeRF[mildenhall2020nerf] is capable of achieving photorealistic scene reconstruction by optimizing a photometric loss between the representation and the images. However, while achieving huge success in the graphics community, the transition of objective function from reprojection error to photometric loss has raised a new challenge to the robotics community: Can we still build consistent scene representations with a moving light source on the robot platform?
In this work, we identify illumination-inconsistency as the main challenge in building a photorealistic scene representation from images taken with a moving light source. That is, as the robot operates in the environment, the effect of moving light source results in images captured of the same region appearing visually different. Handcrafted feature descriptors that serve as backbone in SfM and SLAM [sift, murAcceptedTRO2015, ORBSLAM3_TRO] are designed to be invariant to changes in pixel intensities. Such changes also have limited effects on the optimization pipeline, which uses reprojection errors as objective functions. Hence, the classic SfM methods are generally robust against illumination-inconsistency. However, on the way towards photorealistic scene reconstruction and switching to optimization of the photometric loss, we find that state-of-the-art NeRF methods [mildenhall2020nerf, kerbl3Dgaussians] fail on the images taken in the dark with a moving light source. Although variants such as RawNeRF [mildenhall2022rawnerf] and Relightable 3DG [R3DG2023] claim to be able to handle dark images in RAW format and model the reflection properties of the scene, they still fail to handle the varying illumination.
@@
    DarkGS (DarkGS), a variant of the 3DGS (3DGS) model that builds photorealistic scene representations under poorly illuminated conditions and relights the scene with global illumination, based on COLMAP [colmap] and NeLiS results.
@@ II-A Extroceptive Sensor Calibration on Robots
Autonomous robots are usually equipped with perceptual sensors such as cameras and LiDARs. Taking camera calibration as an example, a calibration target with AprilTag [apriltag] or checkerboard pattern is often used. With observations of the target from different perspectives, the focal length, center of projection, distortion coefficient, and pose of the camera can be estimated [rehder2016extendingkalibr]. Similar target-based approaches have been used to calibrate LiDARs [beltran2022camlidar], radars [radarcameracalib] and acoustic sensors [sonarcalib] for downstream sensor fusion tasks [song2024lirafusion]. Analogous to sensor calibrations mentioned above, in this paper, we propose to use a target consisting of AprilTags [apriltag] and blank space for calibrating the light in a camera-light system, including estimating the transformation between camera’s and light’s coordinate system, RID (RID) and light fall-off curve of the light source.
@@ II-B Light Calibration
Existing methods use various kinds of customized calibration targets: [santo2018light][santo2020light] propose to use a target of AprilTags and pins to infer the position of point light source from the shadows of the pins. Alternatively,  [fastled2021] [Ma:23] propose to use a Lambertian sphere and estimate the light source parameters by learning to reconstruct the sphere. Furthermore, [ShadowBasedLight] shows that the location of the light source can also be recovered from the shadow of a sphere. With a Lambertian plane and customized markers,  [Ma:19] shows that it is possible to calibrate a point light source with a single image. [song2020light] shows the most relevance to our work, which calibrates the pose of a light source and the metric scale given the RID curve. However, the above work either assumes an oversimplified model, e.g. a point light source with inverse square light falloff, in a pure dark environment, or that the RID is given, which makes the method less applicable. In this work, we use a minimum calibration target design that consists of only AprilTags [apriltag] and a Lambertian plane similar to [song2020light]. With NeLiS we show that modeling and learning RID, light falloff and ambient light together with camera-light pose can improve the quality of image reconstruction and make our calibration method more feasible for real-world robot deployments.
@@ II-C NeRF and 3D Gaussian Splatting
Based on the success of SfM [buildrome], NeRF learns to represent scenes using a continuous function, e.g. a MLP (MLP), and can achieve photorealistic novel-view image synthesis [chen2024survey]. Popular variations of NeRF such as Mip-NeRF[Barron2021MipNeRFAM] and Instant-NGP [mueller2022instant] have improved NeRF by rendering accuracy and speed. Recent innovation introduced as 3DGS [kerbl3Dgaussians] has revolutionized this field by replacing the neural network backbone with spatial Gaussians, allowing rendering speed of 100+ fps while maintaining the accuracy and differentiability of the system. For our robotic setup, the problem of varying illumination is attempted by [neuralref, neuralsea] which model the physical property of the scene, e.g. roughness, albedo, and normal, instead of the radiance as a constant. However, these works oversimplify the light model as a Lambertian point light source, co-centered with the camera, making it not applicable on many real robotic platforms. Gaussianshader [jiang2023gaussianshader] and Relightable 3DG [R3DG2023] introduce physical properties into the 3DGS framework but model illumination as a constant that does not change frame-by-frame. According to our experiments, none of the above mentioned methods handles the illumination-inconsistancy issue on a real-robot imaging system. In our proposed DarkGS, by modeling the physical property of the scene and taking advantage of NeLiS, we can not only build a consistent 3DGS from poorly illuminated images, but also relight the scene with global illumination.
@@ II-D Photometric Stereo
Photometric stereo refers to the estimation of the shape of an object with multiple images taken from a single viewpoint with different lighting orientations [photometricstereo]. Early approaches confine this problem to known light source and Lambertian surfaces with uniform albedo. Further work has allowed spatially varying albedo [chakrabarti2016single] and unknown light [photounknownlight]. However, our robotic setup differs from this problem by the fact that our camera and light source always move together as a rigid body. And we choose to tackle this problem within the framework of 3DGS [kerbl3Dgaussians] which allows us to synthesize novel views with photorealistic quality.
@@ III METHODOLOGY
Our proposed workflow for building NeLiS and DarkGS is shown in Fig. 2. We build NeLiS by calibrating the camera-light system at a planar calibration target. By accounting for the calibration image sequence and hand measurement of the light pose, an in-the-loop human operator can initialize the light parameters using our intuitive GUI. Since the calibration target can be localized by AprilTags [apriltag], a shading model can render the image with the given parameters to approximate the calibration image. Then, we run the optimization to minimizes the photometric error of the calibration image sequence and rendered image. After obtaining a calibrated NeLiS, the robot can be deployed in a dark environment, using the collected images to build a DarkGS, finetune the model, and relight the scene.
@@ III-A Shading Model for NeLiS
Given the true size of the calibration target, we can apply PnP (PnP[pnp] to extract $R_{c}^{w}\in SO(3)$ and $t_{c}^{w}\in\mathbb{R}^{3}$ which transform points from the camera coordinate to the world coordinate. The position of the light source is then given by $\mathbf{o}_{l}=R_{c}^{w}t_{l}^{c}+t_{c}^{w}\in\mathbb{R}^{3}$. Since we align the $z$ axis of the light coordinate with the centerline of the light, the direction of the centerline can be denoted by $\boldsymbol{\omega}_{l}=R_{c}^{w}R_{l}^{c}[0,0,1]^{\top}\in\mathbb{R}^{3}$. Both $\mathbf{o}_{l}$ and $\boldsymbol{\omega}_{l}$ are in the world coordinate frame.
@@ III-A1 RID
RID is commonly modeled as a function of the angle between the centerline $\boldsymbol{\omega}_{l}$ of the light and light ray $\boldsymbol{\omega}_{\mathbf{x}}=\mathbf{x}-\mathbf{o}_{l}$. Previous work assume that RID is given [song2020light, Ma:19, Ma:23]. However, this assumption may not hold for in-the-wild robot deployment. Instead, we remove this dependency by learning a neural RID from calibration data:
@@ III-A2 Light Falloff Curve
Inverse square law is widely used to model light falloff, based on the assumption of a point light source. However, when objects are closer to the light source, the inverse square law starts to fail. We choose to model the light falloff in the form of a Lorentzian function of the distance $\lVert\boldsymbol{\omega}_{\mathbf{x}}\rVert_{2}$, as suggested by [ryer1997light]:
@@
Instead of estimating $\tau$ from hand measurement of the light surface [ryer1997light], we designate it to be a learnable parameter.
@@ III-A4 BRDF
Although there are BRDF (BRDF) models capable of handling complex reflection effects, they exceed the scope of this paper. We opt for the Lambertian reflection model, eliminating the need to optimize any parameters in our BRDF. We use $f_{r}(\boldsymbol{\omega}_{\mathbf{x}},\mathbf{n},\mathbf{c})=\text{max}(\boldsymbol{\omega}_{\mathbf{x}}\cdot\mathbf{n},0)\mathbf{c}$ as our BRDF (Lambert cosine law included in BRDF for simplicity, following convention of NeRV [nerv2021] ), giving the rendering equation:
$\hat{L}_{\mathbf{x}}=I_{\mathbf{x}}f_{r}(\boldsymbol{\omega}_{\mathbf{x}},\mathbf{n},\mathbf{c})$ (4)
@@ III-B Building DarkGS
$\hat{L}_{u,v}=\sum_{i\in\mathcal{N}}I_{i}f_{r}(\boldsymbol{\omega}_{i},\mathbf{n}_{i},\mathbf{c}_{i})\alpha_{i}\prod_{j}^{i-1}(1-\alpha_{j})$ (6)
@@ III-B1 Scale Recovery
The framework of 3DGS and its variants are often based on monocular SfM solutions such as COLMAP [buildrome]. While calibration using AprilTags recovers poses in true metric scale with NeLiS, monocular SfM only gives up-to-scale poses for building 3D Gaussians. Here, we introduce a scaling factor $s>0$ as a learnable parameter so that we can obtain the positions with scale $\mathbf{p}_{i}^{\prime}=s\mathbf{p}_{i}$. With captured pixel intensity $L_{u,v}$, the 3D Gaussian can be built by solving the follow optimization problem:
@@ IV-A Experiments Setup
Our imaging system consists of a FLIR machine vision camera (Model Firefly S) and a light source, together mounted on a rigid structure on top of our legged robot, as shown in Fig. 4. The baseline between the camera and the light source is approximately 32 cm, measured by hand. Although the ray-tracing problem can be simplified by placing the light source and camera close enough and modeling them as co-centered [neuralref], in many applications a long baseline between them is a preferred design worth studying. For example, deep-sea robots usually have a long baseline between light and camera to avoid massive light backscattering from the water.
We experiment with three different light sources: a flashlight, a diving light, and a flood light. To build NeLiS, we take $\sim 40$ calibration images from different ranges and perspectives for each light configuration and use $25\%$ of the images as a testing set to evaluate the NeLiS performance. For building DarkGS, we take $50\sim 150$ images for each scene. All images are streamed in RAW format without any tone mapping or white balancing. Parameters optimized in both NeLiS and DarkGS use Adam optimizer [adam] with LieTorch [teed2021tangent] integration.
@@ IV-B Can existing 3DGS methods do the job?
We first investigate whether existing 3DGS methods can reconstruct the scene using the images we collected with our robotic setup in the dark. We experimented with: 1. vanilla 3DGS [kerbl3Dgaussians] which models the scene with constant radiance 2. RawNeRF [mildenhall2022rawnerf] which is developed to reconstruct scene from RAW HDR images 3. Relightable 3DGS [R3DG2023] which models the shading and environmental light map. To make a fair comparison, we apply gamma tone mapping to images for vanilla 3DGS [kerbl3Dgaussians] and Relightable 3DGS [R3DG2023] which brightens the image and smooth the discrepancy between illuminated and under-illuminated areas. We also replace RawNeRF’s backbone MLP-based renderer [mildenhall2022rawnerf] with Gaussian render [kerbl3Dgaussians].
As shown in Fig. 5, we find that none of the existing methods is able to build valid scene reconstructions. We see an excessive amount of artifacts in the center area of the synthesized image. The key reason is that the capability to model varying illumination is missing from existing methods.
Figure 5: None of the existing methods can solve the problem: Results of Vanilla 3DGS [kerbl3Dgaussians], RawNeRF [mildenhall2022rawnerf] Relightable 3DGS [R3DG2023] show heavy artifacts and fail to converge. The key reason for the failures is that the existing method does not model the illumination change.
@@
Figure 6: (a) Light sources on real robots have various RID patterns; (b) Real world measurements of light falloff show that the inverse square law is insufficient to model any of our light sources, but Lorentzian functions [ryer1997light] with learnable parameter $\tau$ fit them well.